An investment will pay $150 at the end of each of the next 3 years, $250 at the end of Year 4, $400 at the end of Year 5, and $550 at the end of Year 6. If other investments of equal risk earn 6% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value: $  

Future value: $  
Answer #1

B D A 1 Solution Present Value of Cashflows Year Cashflow 150 150 150 250 400 550 PVE @ 6% 0.943 0.890 0.840 0.792 0.747 0.70
Answer #2

q Present Value Annuity: An investment will pay $150 at the end of each of the next 5 years, If other investments of equal risk earn 11% annually, what is its present value?

source: Chapter 5 Time Value Of Money Page : 148
answered by: musa
Answer #3

SOLUTION :


Discount rate, r = 6% = 0.06 

=> 1 = r = 1.06


PV 

= 150/1.06+150/1.06^2+150/1.06^3+250/1.06^4+400/1.06^5+550/1.06^6

= 1285.61 ($) (ANSWER).


FV 

= PV*1.06^6 

= 1285.61*1.06^6

= 1823.66 ($) (ANSWER)

answered by: Tulsiram Garg

> In the 2nd line, please read " 1 = r = 1.06 " as " 1 + r = 1.06 "

Tulsiram Garg Thu, Oct 14, 2021 2:52 AM